OpenStudy (emilyjones284):

Please help ...medal and fan (: A set of data has a normal distribution with a mean of 5.1 and a standard deviation of 0.9. Find the percent of data between 4.2 and 5.1.

OpenStudy (nurali):

You need to first calculate z scores for the lower limit of your interval: z = (x - µ) / σ = (4.2 - 5.1) / .9 = -.9 / .9 = -1.0 Then you need to look up z = -1 on a normal distribution table. It shows that the proportion of cases falling between the mean and a z of -1 is .3413 or 34.13%

OpenStudy (emilyjones284):

Okay got it thanks (:

OpenStudy (nurali):

My Pleasure.